f([x^2 + 1]/x)=x^2 + 1/x^2,x is not equal to 0
then f(x)=?

Grade:12th pass

1 Answers

Khimraj
6 years ago
simply put x2 = t-1
f(t) = t/(t-1)
Now put t = x
f(x) = x/(x-1).
Hope it clears......................................